OpenStudy (anonymous):

A bus moves away from a bus stop in a straight line. It moves 10 meters by the end of the 1st second, 20 meters by the end of the 2nd second, and 30 meters by the end of the 3rd second. How many meters away is the bus from the bus stop after 10 seconds? a. 100 c. 120 b. 110 d. 190

OpenStudy (popitree):

in each second it moves 10m

OpenStudy (popitree):

so figure out where it will be after 10 sec

OpenStudy (anonymous):

speed = 10m/s therefore distance=10*10=100m